OpenStudy (anonymous):

Fool's problem of the day, Find the value of \( x^4 +9x^3+35x^2-x+4\) for \( x = -5+2\sqrt{-4}\)

OpenStudy (anonymous):

Ishaan I am expecting an answer of you :)

OpenStudy (anonymous):

umm yeah, I am trying to find an easier, smart and quick way to do it.

OpenStudy (anonymous):

well of-course, the standard substation is not feasible without electronic aid.

OpenStudy (anonymous):

this is embarassing it seems so easy

OpenStudy (anonymous):

I agree ishaan :D

OpenStudy (anonymous):

Here comes mighty asnaseer!

OpenStudy (asnaseer):

isn't this just a simple substitution of \(x=-5+4i\) into the above equation?

OpenStudy (anonymous):

Yes asnaseer but that would be messy/tedious isn't ?

OpenStudy (asnaseer):

the equation could be rearranged to:\[x(x(x(x+9)+35)-1)+4\]to simplify the calculations

OpenStudy (asnaseer):

or maybe find some rearrangement to get terms involving:\[x(x-8i)=(-5+4i)(-5-4i)=25+16=41\]

OpenStudy (anonymous):

You did it again asnaseer! :)

OpenStudy (asnaseer):

I did? I know the answer but I am trying to see how I can make use of my last idea?

OpenStudy (anonymous):

160 but this was so embarassing for me :-/

OpenStudy (asnaseer):

-160

OpenStudy (anonymous):

I like this one \( x(x(x(x+9)+35)-1)+4 \)

OpenStudy (asnaseer):

thats the one I used.

OpenStudy (asnaseer):

I thought maybe \(x(x-8i)\) might be useable somehow but can't see how.

OpenStudy (anonymous):

I can't either.

OpenStudy (asnaseer):

:) phew! - so you ARE mortal :D

OpenStudy (anonymous):

lol, you may never know :P

OpenStudy (anonymous):

Today is a bad day for me, example : http://openstudy.com/users/foolformath#/updates/4f367ecde4b0fc0c1a0ce793